# EgoTraj-Bench: Towards Robust Trajectory Prediction Under Ego-view Noisy Observations

### v1

@@ Abstract
Reliable trajectory prediction from an ego-centric perspective is crucial for robotic navigation in human-centric environments. However, existing methods typically assume idealized observation histories, failing to account for the perceptual artifacts inherent in first-person vision, such as occlusions, ID switches, and tracking drift. This discrepancy between training assumptions and deployment reality severely limits model robustness. To bridge this gap, we introduce EgoTraj-Bench, the first real-world benchmark that grounds noisy, first-person visual histories in clean, bird’s-eye-view future trajectories, enabling robust learning under realistic perceptual constraints. Building on this benchmark, we propose BiFlow, a dual-stream flow matching model that concurrently denoises historical observations and forecasts future motion by leveraging a shared latent representation. To better model agent intent, BiFlow incorporates our EgoAnchor mechanism, which conditions the prediction decoder on distilled historical features via feature modulation. Extensive experiments show that BiFlow achieves state-of-the-art performance, reducing minADE and minFDE by 10–15% on average and demonstrating superior robustness. We anticipate that our benchmark and model will provide a critical foundation for developing trajectory forecasting systems truly resilient to the challenges of real-world, ego-centric perception.
@@ I Introduction
Pedestrian trajectory prediction[1, 2], aiming to estimate the multimodal future paths of agents in dynamic environments, serves as a foundation for safe, socially compliant motion planning in autonomous systems such as mobile robots, intelligent prosthetics, and service vehicles[3, 4, 5, 6, 7, 8]. Although extensively studied, most existing methods are developed and evaluated under idealized bird’s-eye view (BEV) settings with globally consistent observations and flawless agent tracking[2]. However, these conditions rarely hold in real-world deployment. Autonomous agents, such as mobile robots, typically perceive the environment through front-facing cameras, where observations are inherently incomplete and noisy as illustrated in Fig. 1: pedestrians may be occluded, enter or exit the field of view (FOV), or suffer from physical sensing errors such as perspective distortion. These imperfections substantially violate the idealized historical assumptions in BEV settings. Therefore, trajectory prediction under ego-centric (first-person view, FPV) noisy observations is essential for enabling robust deployment in real-world scenarios.
Figure 1: Illustration of key challenges under ego-view observations. Top row: Occlusion. In the first-person view (a), only one pedestrian (green box) is visible due to occlusion; the corresponding BEV (b) shows two additional agents (pink and yellow) who are behind the green pedestrian. Dashed lines indicate trajectories visible in BEV but not in FPV. Bottom row: ID Switch and Perspective Distortion. Two pedestrians (yellow and pink) cross paths, causing an ID swap in the FPV tracking output (a). Additionally, individuals near the image corners suffer from significant perspective distortion, making accurate localization challenging.
Some prior work [9, 6, 10] predicts trajectories from ego-centric input, typically predicting future positions in image space, e.g., bounding boxes or keypoints. Although operating from an ego-centric view, these methods lack spatial reasoning in real-world space and assume idealized tracking results in image space, leaving unresolved the modeling of fine-grained interaction in real-world trajectory prediction. In contrast, methods[11, 12] that predict trajectories in global metric spaces (e.g., world coordinates) enable precise spatial reasoning about proximity, collision risk, and social norms, and we therefore focus on the latter paradigm.
@@
Figure 2: EgoTraj-Bench Overview: Left Synchronized BEV and FPV videos are obtained from the dataset. Blue box marks a temporally aligned frame. Mid Clean past and future trajectories are extracted from BEV annotations as ground truth, while noisy historical observations are projected from FPV videos. Right The noisy ego-view histories are paired with ground truth, enabling robust evaluation under realistic ego-centric conditions. A mask is also generated based on history visibility.
@@ II-A Pedestrian Trajectory Prediction
Various models have been proposed for pedestrian trajectory prediction. Social-LSTM[1] pioneers agent interaction modeling via pooled LSTM states. Social GAN[2] and AC-VRNN[18] introduce generative frameworks to capture multimodal futures. More recently, Transformer-based architectures like TUTR[19] leverage self-attention for long-range spatiotemporal reasoning. Denoising models, particularly diffusion models and flow-based models[20, 21, 22], effectively modeling complex multimodal trajectory distributions via iterative denoising. Beyond architecture, several methods incorporate auxiliary inputs, e.g, goal estimates[23, 12] or scene graphs[11], to enrich contextual cues. However, these typically assume clean, globally observed trajectories, an assumption violated in real-world ego-centric settings. While some studies attempt to enhance robustness via random history masking[24, 25, 26], such artificial missingness fails to reflect the complex and structured nature of real ego-centric perception. This gap motivates the need for trajectory prediction approaches under realistic ego-centric perturbations.
@@ II-B Trajectory Prediction under Ego-centric Noise
Methodologically, T2FPV’s CoFE module denoises historical perturbations using clean history as supervision[14], but only corrects missing positions when constructing the input of prediction model. This hybrid representation, combining uncorrected observed positions with corrected ones, risks spatial-temporal discontinuities at segment boundaries, potentially causing artificial motion jumps during prediction. To this end, we propose a unified framework avoiding such patchwork reconstruction with the first real-world benchmark that captures both noisy perception inputs and metric-accurate ground truth.
II-C Diffusion and Flow Models
Inspired by non-equilibrium statistical physics, diffusion models balance tractability and flexibility by gradually corrupting data through a forward diffusion process and learning to reverse it using stochastic dynamics[30]. They are typically formulated as stochastic differential equations (SDEs) that map a noise distribution to the data distribution, which enables high-fidelity generation while suffering from slow sampling. To accelerate inference, recent work converts the SDE into a deterministic ordinary differential equation, enabling faster, deterministic generation. This shift motivates the use of flow matching (FM)[31], which directly learns the velocity field guiding future positions from noise to data. FM has proven effective in stabilizing training and improving sample quality across 3D, video, and graph generation[32, 33, 34, 35, 36].
In trajectory prediction, MID[20] first applies diffusion to predict future trajectory, while Leapfrog[21] introduces a deterministic latent initializer to speed up sampling. However, both sample futures independently per agent, leading to spatially redundant trajectories. MoFlow[22] uses FM to jointly model futures conditioned on past motion, promoting learning of diverse future trajectories. Thus, we adopt FM to generate future trajectories from partial and noisy ego-centric inputs, leveraging its ability to condition on corrupted histories and support diverse predictions.
@@ III EgoTraj-Bench: Ego-view Trajectory Prediction Benchmark
Dataset Fold Duration Label Freq # Traj Ego-noise Involved BEV Future FPV Noisy Rate History MSE
@@
WildTrack[37] 3 2 313
THöR[38] 60 100 600
T2FPV–ETH[14] ETH 9 15 181 0.44 5.66
Hotel 13 15 1,053 0.51 4.55
Zara1 6 2.5 5,939 0.28 1.23
Zara2 7 2.5 17,608 0.32 2.50
Univ 3.5 2.5 24,334 0.45 2.47
EgoTraj-TBD 210 30 36,947 0.37 0.66
  • Further details, such as density statistics, are available in TBD[17].
@@ III-A Real-world Dataset Creation
- Trajectory Extraction. To ensure perception noise reflects real deployment conditions, we extract pedestrian trajectories from raw FPV video using YOLOv8[39] for detection, selected for its robustness in crowded scenes, and BotSort[40] for tracking, which fuses motion and appearance cues to mitigate ID switches. Visibility is quantified per frame via segmentation masks from YOLOv8-seg. All hyperparameters are tuned for scenes. The final 2D bounding box bottom centers are back-projected to the ground plane using calibrated intrinsics and synchronized ego-motion from the TBD dataset, yielding BEV-space trajectories that retain realistic noise such as occlusion, ID instability, and localization error. For the supervision of clean past and future trajectories, we use the BEV annotations provided in the TBD dataset, benefiting from occlusion-resilient multi-view coverage and semi-automated human verification.
@@ III-B Statistics and Analysis
EgoTraj-TBD covers two key challenge types: (1) inherent perceptual artifacts (e.g., occlusion, FOV truncation, ID switching); and (2) physical sensor artifacts (e.g., lens distortion, projection error, ego-motion drift), important to real-world systems. Our dataset features a moderate noise rate (average probability of being marked as invisible) and lower historical alignment error (distance between estimated and ground-truth history), which could be explained by higher-fidelity ground truth and a more noise-aware processing pipeline as mentioned in Sec. III-A. During learning, the dataset is chronologically split into train, validation, and test sets (70%-10%-20%) to ensure temporal coherence and avoid data leakage.
Figure 3: Overview of our BiFlow. The input consists of a noisy historical trajectory $\tilde{X}$ and its corresponding visibility mask $m$. During training, the model is supervised with clean ground-truth past $X$ and future $Y$ trajectories to jointly learn reconstruction and prediction. At inference, only the noisy history and mask are used as input to predict the future trajectory $\hat{Y}$.
@@ III-C Evaluation Framework and Findings
- Empirical Findings. As shown in Table II, our benchmark reveals a significant gap: all BEV-trained state-of-the-art models exhibit substantial degradation under ego-view perception noise. Particularly, on the widely adopted ETH-UCY datasets, the minADE@20 rises to 0.67m, compared to approximately 0.20m under clean historical trajectories[19, 22]. This highlights a critical limitation that existing methods are highly sensitive to perception artifacts prevalent in ego-view data. These findings underscore the need for noise-aware modeling in trajectory prediction.
@@ IV-A Problem Formulation
We consider the task of multi-agent trajectory prediction under ego-centric observation noise. Let $A$ denote the total number of agents in the scene. The input consists of: (1) the observed history $T_{p}$ steps of all agents $\tilde{X}\in\mathbb{R}^{A\times 2T_{p}}$, structured as $\tilde{X}=[\tilde{X}_{\text{other}};\tilde{X}_{\text{ego}}]$, where $\tilde{X}_{\text{other}}\in\mathbb{R}^{(A-1)\times 2T_{p}}$ contains noisy positions of non-ego agents observed from FPV video, and $\tilde{X}_{\text{ego}}\in\mathbb{R}^{2T_{p}}$ contains clean, fully observable positions of the ego-agent from robot odometry; (2) a binary visibility mask $m\in\{0,1\}^{A\times T_{p}}$, structured as $m=[m_{\text{other}};\mathbf{1}_{T_{p}}]$, where $m_{\text{other}}\in\{0,1\}^{(A-1)\times T_{p}}$ indicates visibility of non-ego agents, and $\mathbf{1}_{T_{p}}$ denotes full observability of the ego-agent across all $T_{p}$ steps. The goal is to predict the future trajectories $Y\in\mathbb{R}^{A\times 2T_{f}}$ of all agents over the next $T_{f}$ time steps.
@@ IV-B Overview
Motivated by our benchmark findings that ego-view noise severely degrades the trajectory prediction performance of existing methods, we propose BiFlow, a dual-stream framework that jointly reconstructs clean agent history trajectories and predicts their future trajectories.
@@ IV-C Contextual Encoder: Modeling Noisy Social Dynamics
To capture the noisy dynamics and social interactions of agents’ past trajectories, we design an agent-aware contextual encoder using the Transformer architecture. The input is embedded with historical trajectory data $\tilde{X}$ and visibility masks $m$:
$H_{0}=\mathrm{MLP}\left(\tilde{X}\oplus m\right).$ (4)
We first apply multi-head self-attention (MHSA) to learn features that model the semantics of agents’ social interactions:
$H_{\text{soc}}=\mathrm{MHSA}(Q=H_{0},\,K=H_{0},\,V=H_{0}).$ (5)
A second MHSA layer is then applied to refine the learned representation into coherent global scene representations as follows:
@@
where $\mathrm{PE}_{A}$ is a learnable positional encoding based on agent identities. The output $H_{\text{enc}}\in\mathbb{R}^{A\times D}$ serves as the shared latent representation for both reconstruction and prediction streams, where D is the feature dimension.
@@ IV-D EgoAnchor: Intent Prior Distillation
where $Ach_{\text{agent}}\in\mathbb{R}^{A\times D}$ is used to capture agent-level motion tendencies, and $Ach_{\text{scene}}\in\mathbb{R}^{D}$ summarizes global context. We use layer normalization to handle crowd size, visibility, and camera noise variations in $H_{enc}$, making the output anchor comparable across agents and scenes.
@@ IV-E Dual Decoder with Multi-Candidate Prediction
The history reconstruction stream mirrors identical structure but sets $\beta_{Ach}$ and $\gamma_{Ach}$ to 0, enabling pure reconstruction learning. Its training objective follows the same form as that of the prediction branch:
$L_{recon}$ $=\mathbb{E}_{t,\,X_{1},\,X_{t}}\!\left[\bigl\|\hat{X}^{1:K}-X_{1}\bigr\|_{2}^{2}+\mathrm{CE}(c_{x}^{1:K},j^{*}_{x})\right],$ (15)
$j^{*}_{x}$ $=\arg\min_{j}\|\hat{X}^{j}-X_{1}\|_{2}^{2}$ (16)
During inference, the history reconstruction decoder is no longer activated. Instead, we use only the noisy input trajectory $\tilde{X}$, which is processed through the shared contextual encoder and the prediction decoder and conditioned on intent priors from the EgoAnchor module, to forecast future trajectories.

### v2

@@ Abstract
Reliable trajectory prediction from an ego-centric perspective is crucial for robotic navigation in human-centric environments. However, existing methods typically assume noiseless observation histories, failing to account for the perceptual artifacts inherent in first-person vision, such as occlusions, ID switches, and tracking drift. This discrepancy between training assumptions and deployment reality severely limits model robustness. To bridge this gap, we introduce EgoTraj-Bench, built upon TBD dataset, which is the first real-world benchmark that aligns noisy, first-person visual histories with clean, bird’s-eye-view future trajectories, enabling robust learning under realistic perceptual constraints. Building on this benchmark, we propose BiFlow, a dual-stream flow matching model that concurrently denoises historical observations and forecasts future motion. To better model agent intent, BiFlow incorporates our EgoAnchor mechanism, which conditions the prediction decoder on distilled historical features via feature modulation. Extensive experiments show that BiFlow achieves state-of-the-art performance, reducing minADE and minFDE by 10–15% on average and demonstrating superior robustness. We anticipate that our benchmark and model will provide a critical foundation for robust real-world ego-centric trajectory prediction. The benchmark library is available at: https://github.com/zoeyliu1999/EgoTraj-Bench.
@@ I Introduction
Pedestrian trajectory prediction[1, 2], aiming to estimate the multimodal future paths of agents in dynamic environments, serves as a foundation for safe, socially compliant motion planning in autonomous systems such as mobile robots, intelligent prosthetics, and service vehicles[3, 4, 5, 6, 7]. Although extensively studied, most existing methods are developed and evaluated under idealized bird’s-eye view (BEV) settings with globally consistent observations and flawless agent tracking[2]. However, these conditions rarely hold in real-world deployment. Autonomous agents, such as mobile robots, typically perceive the environment through front-facing cameras, where observations are inherently incomplete and noisy as illustrated in Fig. 1: pedestrians may be occluded, enter or exit the field of view (FOV), or suffer from physical sensing errors such as perspective distortion. These imperfections substantially violate the idealized historical assumptions in BEV settings. Therefore, trajectory prediction under ego-centric (first-person view, FPV) noisy observations is essential for enabling robust deployment in real-world scenarios.
Figure 1: Illustration of key challenges under ego-view observations. Top row: Occlusion. In the first-person view (a), only one pedestrian (green box) is visible; the corresponding BEV (b) shows two additional agents (pink and yellow) behind the green pedestrian. Dashed lines indicate trajectories visible in BEV but not in FPV. Bottom row: ID Switch and Perspective Distortion. Two pedestrians (yellow and pink) cross paths, causing an ID swap in the FPV tracking output (a). Additionally, individuals near the image corners suffer significant perspective distortion, making accurate localization challenging.
Some prior work [8, 6, 9] predicts trajectories from ego-centric input, typically predicting future positions in image space, e.g., bounding boxes or keypoints. Despite operating from an ego-centric view, these methods lack spatial reasoning in real-world space, assume idealized tracking results in image space, and thus leave the modeling of fine-grained interactions unresolved in real-world trajectory prediction. In contrast, methods[10, 11] that predict trajectories in global metric spaces (e.g., world coordinates) enable precise spatial reasoning about proximity, collision risk, and social norms, and we therefore focus on the latter paradigm.
@@
Figure 2: EgoTraj-Bench Overview: Left contrasts the same scene in BEV vs FPV, where limited visibility and viewpoint changes introduce noisy histories. Mid shows sample construction: synchronized BEV–FPV clips provide noisy FPV-derived past trajectories as inputs and clean metric BEV trajectories as supervision, with a history validity mask. Right illustrates what the benchmark records and the key noise patterns that drive robustness gaps beyond idealized BEV settings.
@@ II-A Pedestrian Trajectory Prediction
Various models have been proposed for pedestrian trajectory prediction. Social-LSTM[1] pioneers agent interaction modeling via pooled LSTM states. Social GAN[2] and AC-VRNN[17] introduce generative frameworks to capture multimodal futures. More recently, Transformer-based architectures like TUTR[18] leverage self-attention for long-range spatiotemporal reasoning. Denoising models, particularly diffusion models and flow-based generative models[19], further improve the modeling of complex multimodal trajectory distributions via iterative refinement. MID[20] first applies diffusion to predict future trajectory, while Leapfrog[21] introduces a deterministic latent initializer to speed up sampling. However, both sample futures independently per agent, leading to spatially redundant trajectories. MoFlow[22] uses flow matching (FM) to jointly model futures conditioned on past motion, promoting learning of diverse future trajectories. These advances motivate adopting FM to generate future trajectories from partial and noisy ego-centric inputs, leveraging its ability to condition on corrupted histories and support diverse predictions.
Beyond architecture, several methods incorporate auxiliary inputs, e.g., goal estimates[23, 11] or scene graphs[10], to enrich contextual cues. However, these typically assume clean, globally observed trajectories, an assumption violated in real-world ego-centric settings. While some studies attempt to enhance robustness via random history masking[24, 25, 26], such artificial missingness fails to reflect the complex and structured nature of real ego-centric perception. This gap motivates the need for trajectory prediction approaches under realistic ego-centric perturbations.
@@ II-B Trajectory Prediction under Ego-centric Noise
Methodologically, T2FPV’s CoFE module denoises historical perturbations using clean history as supervision[13], but only corrects missing positions when constructing the input of prediction model. This hybrid representation, combining uncorrected observed positions with corrected ones, risks spatial-temporal discontinuities at segment boundaries, potentially causing artificial motion jumps during prediction. To this end, this motivates a unified framework that avoids such patchwork reconstruction, together with a real-world benchmark capturing both noisy perception inputs and metric-accurate ground truth.
@@ III EgoTraj-Bench: Ego-view Trajectory Prediction Benchmark
Dataset Fold Duration Label Freq # Traj Ego-noise Involved Agent Scale FPV Noisy Rate History MSE
@@
WildTrack[30] 3 2 313 pedestrian
THöR[31] 60 100 600 pedestrian
nuScenes[27] 330 12 - ✗ vehicle
T2FPV–ETH [13] ETH 9 15 181 pedestrian 0.44 5.66
Hotel 13 15 1,053 pedestrian 0.51 4.55
Zara1 6 2.5 5,939 pedestrian 0.28 1.23
Zara2 7 2.5 17,608 pedestrian 0.32 2.50
Univ 3.5 2.5 24,334 pedestrian 0.45 2.47
EgoTraj-TBD 210 30 36,947 pedestrian 0.37 0.66
  • $\dagger,*$
    FPV noisy rate indicates the average probability of being marked as invalid. History MSE indicates the distance between estimated and ground-truth history. Further details, such as density statistics, are available in TBD[16].
@@ III-A Real-world Dataset Creation
- Trajectory Extraction. To ensure perception noise reflects real deployment conditions, we extract pedestrian trajectories from raw FPV video using YOLOv8[32] for detection, selected for its robustness in crowded scenes, and BotSort[33] for tracking, which fuses motion and appearance cues to mitigate ID switches. Visibility is quantified per frame via segmentation masks from YOLOv8-seg. All hyperparameters are tuned for different scenes. The final 2D bounding box bottom centers are back-projected to the ground plane using calibrated intrinsics and synchronized ego-motion from the TBD dataset, yielding BEV-space trajectories that retain realistic noise such as occlusion, ID instability, and localization error. For the supervision of clean past and future trajectories, we use the BEV annotations provided in the TBD dataset, benefiting from occlusion-resilient multi-view coverage and semi-automated human verification.
@@ III-B Statistics and Analysis
EgoTraj-TBD covers two key challenge types: (1) inherent perceptual artifacts (e.g., occlusion, FOV truncation, ID switching); and (2) physical sensor artifacts (e.g., lens distortion, projection error, ego-motion drift), important to real-world systems. EgoTraj-Bench also focuses on pedestrian-scale agents in human-centric environments, whose surroundings and motion patterns differ fundamentally from vehicle-centric benchmarks such as nuScenes [27]. As reported in Table I, the dataset features a moderate noise rate and lower historical alignment error, which could be explained by higher-fidelity ground truth and a more noise-aware processing pipeline as mentioned in Sec. III-A. During learning, the dataset is chronologically split into train, validation, and test sets (70%-10%-20%) to ensure temporal coherence and avoid data leakage.
Figure 3: Overview of our BiFlow. The input consists of a noisy historical trajectory $\tilde{X}$ and its corresponding validity mask $m$. During training, the model is supervised with clean ground-truth past $X$ and future $Y$ trajectories to jointly learn reconstruction and prediction. At inference, only the noisy history and mask are used as input to predict the future trajectory $\hat{Y}$.
@@ III-C Evaluation Framework and Findings
- Empirical Findings. As shown in Table II, our benchmark reveals a significant gap: all BEV-trained state-of-the-art models exhibit substantial degradation under ego-view perception noise. Particularly, on the widely adopted ETH-UCY datasets, the minADE@20 rises to 0.67m, compared to approximately 0.20m under clean historical trajectories[22]. This highlights a critical limitation that existing methods are highly sensitive to perception artifacts prevalent in ego-view data. These findings underscore the need for noise-aware modeling in trajectory prediction.
@@ IV-A Problem Formulation
We consider the task of multi-agent trajectory prediction under ego-centric observation noise. Let $A$ denote the total number of agents in the scene. The input consists of: (1) the observed history $T_{p}$ steps of all agents $\tilde{X}\in\mathbb{R}^{A\times 2T_{p}}$, structured as $\tilde{X}=[\tilde{X}_{\text{other}};\tilde{X}_{\text{ego}}]$, where $\tilde{X}_{\text{other}}\in\mathbb{R}^{(A-1)\times 2T_{p}}$ contains noisy positions of non-ego agents observed from FPV video, and $\tilde{X}_{\text{ego}}\in\mathbb{R}^{2T_{p}}$ contains clean, fully observable positions of the ego-agent from robot odometry; (2) a binary validity mask $m\in\{0,1\}^{A\times T_{p}}$, structured as $m=[m_{\text{other}};\mathbf{1}_{T_{p}}]$, where $m_{\text{other}}\in\{0,1\}^{(A-1)\times T_{p}}$ indicates validity of non-ego agents, and $\mathbf{1}_{T_{p}}$ denotes full observability of the ego-agent across all $T_{p}$ steps. The goal is to predict the future trajectories $Y\in\mathbb{R}^{A\times 2T_{f}}$ of all agents over the next $T_{f}$ time steps.
@@ IV-B Overview
Motivated by our benchmark findings that ego-view noise severely degrades the trajectory prediction performance of existing methods, and by the strong performance and efficient conditional generation of recent flow-matching approaches, we propose BiFlow, a dual-stream framework that jointly reconstructs clean agent history trajectories and predicts their future trajectories.
@@ IV-C Contextual Encoder: Modeling Noisy Social Dynamics
To capture the noisy dynamics and social interactions of agents’ past trajectories, we design an agent-aware contextual encoder using the Transformer architecture.
Given the noisy historical trajectory data $\tilde{X}$ and validity masks $m$, we first apply multi-head self-attention (MHSA) to model inter-agent social dependencies. We then add a learnable agent-identity encoding $\mathrm{PE}_{A}$ and apply a second MHSA layer to aggregate a coherent scene-level representation:
$H_{0}$ $=\mathrm{MLP}\left(\tilde{X}\oplus m\right),$ (4)
$H_{\text{soc}}$ $=\mathrm{MHSA}(Q=H_{0},\,K=H_{0},\,V=H_{0}),$ (5)
@@
The output $H_{\text{enc}}\in\mathbb{R}^{A\times D}$ serves as the shared latent representation for both reconstruction and prediction streams, where D is the feature dimension.
@@
  • $\dagger$
    ETH/HOTEL/ZARA1/ZARA2/UNIV are the five folds from T2FPV-ETH.
@@ IV-D EgoAnchor: Intent Prior Distillation
where $Ach_{\text{agent}}\in\mathbb{R}^{A\times D}$ is used to capture agent-level motion tendencies, and $Ach_{\text{scene}}\in\mathbb{R}^{D}$ summarizes global context. We use layer normalization to handle crowd size, validity, and camera noise variations in $H_{enc}$, making the output anchor comparable across agents and scenes.
@@ IV-E Dual Decoder with Multi-Candidate Prediction
The history reconstruction stream mirrors identical structure but sets $\beta_{Ach}$ and $\gamma_{Ach}$ to 0, enabling pure reconstruction learning. During inference, the history reconstruction decoder is no longer activated. Instead, we use only the noisy input trajectory $\tilde{X}$, which is processed through the shared contextual encoder and the prediction decoder and conditioned on intent priors from the EgoAnchor module, to forecast future trajectories.